Find two consecutive even integers x and x+2 with a product of 80
Answer by checkley77(12569)

x(x+2)=80
x^2+2x-80=0
(x+10)(x-8)=0
x+10=0
x=-10 ans.
-10+2=-8 ans
OR:
x-8=0
x=8 ans.
8+2=10 ans.